# Connect the Dots: Knowledge Graph–Guided Crawler Attack on Retrieval-Augmented Generation Systems

### v1

@@ 8 Conclusion
In this work, we formalized the critical privacy risk of data extraction from RAG systems as an adaptive stochastic coverage problem. We introduced RAGCrawler, a novel attack framework that builds a knowledge graph to track revealed information. Our comprehensive experiments demonstrated that RAGCrawler significantly outperforms all baselines, achieving up to 84.4% corpus coverage within a fixed query budget. Furthermore, it shows remarkable robustness against advanced RAG that employs query rewriting and multi-query retrieval. Our findings expose a fundamental vulnerability in current RAG architectures, underscoring the urgent need for robust safeguards to protect private knowledge bases.
9 Appendix
9.1 Sample Distribution Validation
Energy Distance quantifies the difference between two distributions using Euclidean distances, and is well-suited for high-dimensional or non-Gaussian data. A smaller energy distance with a large permutation test $p$-value suggests no statistically significant difference between the sampled and full sets. As shown in Table VII, all four corpora yield low distances (around 0.03) and high $p$-values (>0.5>0.5), indicating strong alignment between samples and full sets. C2ST (Classifier Two-Sample Test) reframes distribution comparison as a binary classification task . If a classifier cannot distinguish between samples and the full corpus (AUC $\approx 0.5$), the two distributions are considered similar. In Table VII, all AUC scores fall within the 0.46–0.51 range, reinforcing that no meaningful distinction can be learned between the subsets and their corresponding full corpora.
TABLE VII: Semantic distribution similarity between each full corpus and its 1,000-document sampled subset, measured by Energy Distance (with permutation test $p$-value) and the Classifier Two-Sample Test (C2ST, using AUC). A small Energy Distance with a high $p$-value and an AUC near 0.5 indicate distributional similarity.
9.2 Proof of Theorems
@@
(Adaptive Monotonicity) The coverage function $f(S,\Phi)$ is monotone with respect to the query set $S$. Issuing an additional query $q$ can only increase or preserve the number of unique documents retrieved. That is, for any realization of $\Phi$,
$f(S\cup\{q\},\Phi)\geq f(S,\Phi).$
Taking the expectation conditioned on $\psi$ preserves this inequality, so $\Delta(q\mid\psi)\geq 0$.
(Adaptive Submodularity) Suppose $\psi^{\prime}$ extends $\psi$, meaning that the attacker has issued more queries and observed more information. Let $S$ and $S^{\prime}$ denote the corresponding sets of issued queries, where $S\subseteq S^{\prime}$. Since more of the corpus has potentially been covered under $S^{\prime}$, the additional gain from issuing a new query $q$ can only decrease. Formally, for any realization of $\Phi$, we have
$f(S\cup\{q\},\Phi)-f(S,\Phi)\geq f(S^{\prime}\cup\{q\},\Phi)-f(S^{\prime},\Phi).$
Taking expectation conditioned on $\psi$ and $\psi^{\prime}$ respectively, we obtain
$\Delta(q\mid\psi)\geq\Delta(q\mid\psi^{\prime}).$
9.3 Additional Ablations
Figure 11: Hyperparameter Choices of RAGCrawler.
Hyperparameter Choices. We examine two key hyperparameters in RAGCrawler: the UCB exploration coefficient $c$ used by the Strategy Scheduler for query selection, and the similarity threshold $\tau$ used by the Query Generator to filter high-similarity queries (thereby controlling early stopping). As shown in Fig. 11, a moderate UCB coefficient of $c=0.5$ delivers the best performance by balancing exploitation and exploration. Increasing or decreasing $c$ from this value noticeably degrades performance, underscoring that $c=0.5$ is near-optimal for a proper exploration–exploitation trade-off. Similarly, the threshold $\tau$ exhibits a clear sweet spot. Setting $\tau$ too low often causes premature termination. On the other hand, an overly high threshold produces queries that are not sufficiently distinct from one another, which limits exploration efficiency. Our chosen $\tau=0.8$ navigates between these extremes.
9.4 Coverage Rate Curves
@@
9.5 Prompts for Experimental Stages
@@
RAG System Prompt You are an assistant for question-answering tasks. Use the following pieces of retrieved context to answer the question. If you don’t know the answer, say that you don’t know. Context: [context] Query::[query] Answer:
Query Rewriting Prompt in RAG System You are an helpful assistant. Your task is to rewrite the given user question to clarify the user’s intent and remove any adversarial intent. Provide directly output the rewritten question without any other text. Original Question: [query] Answer:
Multi-query Prompt in RAG System You are an helpful assistant. Your task is to generate 3 different versions of the given user question to retrieve relevant documents from a vector database. By generating multiple perspectives on the user question, your goal is to help the user overcome some of the limitations of the distance-based similarity search. Provide directly output these alternative questions separated by newlines without any other text. Original Question: [query] Answer:
9.6 Cost of the attack
We estimate the cost of executing RAGCrawler under realistic conditions. As reported in Table VIII, the attacker processes approximately 6–10 million tokens per dataset. Based on Doubao-1.5-lite-32K’s API pricing (\$0.042 per million input tokens and \$0.084 per million output tokens), the total cost per attack is remarkably low, i.e., roughly \$0.33 to \$0.53 per dataset. Moreover, as shown in Sec. 6.4, substituting a freely available open-source model such as Qwen-2.5-7B does not degrade the performance of RAGCrawler, enabling the attack to be executed at virtually zero cost, aside from any fees for the RAG service and GPU hosting.
TABLE VIII: Token usage and estimated cost per corpus for KG-Constructor (KG-C) and Query Generator (QG) using Doubao-1.5-lite-32k. “InTok” and “OutTok” denote input and output token counts. Each attack costs under \$1.
9.7 Additional Experiments Results

### v2

@@ 9 Conclusion
We investigated knowledge-base stealing in RAG systems, a growing threat to the intellectual property of the retrieval corpus. We formalized the attacker’s goal as RAG Crawling, an instance of ASCP that targets global corpus coverage under a fixed query budget, and we developed RAGCrawler, a knowledge graph-guided framework that operationalizes this objective in practical black-box settings. Extensive experiments show that RAGCrawler achieves substantially higher coverage than prior attacks (up to 84.4% within a budget of 1,000 queries) and remains effective under query rewriting and multi-query retrieval, highlighting urgent gaps in protecting the knowledge assets of RAG deployments.
Appendix
Appendix A Proof of Theorems
Notation (issued-query set).
A retrieval-level partial realization $\psi^{R}$ records the issued queries and their (retrieval) outcome sets. We write $\mathrm{dom}(\psi^{R})\subseteq Q$ for the set of queries appearing in $\psi^{R}$ (i.e., issued so far), and use the shorthand
$S(\psi^{R})\ :=\ \mathrm{dom}(\psi^{R}).$
In particular, the $S_{t-1}$ in Definition 1 is exactly $S(\psi^{R}_{t-1})$. We also define the union of revealed retrieved items in $\psi^{R}$ as
$R(\psi^{R})\ :=\ \bigcup_{(q,o)\in\psi^{R}}o.$
Lemma A.1 (RAG $\Rightarrow$ independence of unissued outcomes). We consider standard RAG deployment: each query is processed independently, and any internal randomness (e.g., query rewriting sampling or multi-query generation sampling) is freshly drawn per query and independent across queries. Formally, let $\Phi=(\xi_{q})_{q\in Q}$ where $\{\xi_{q}\}$ are mutually independent, and assume $O_{\Phi}(q)$ is determined only by $(q,\xi_{q})$. Then for any retrieval-level partial realization $\psi^{R}$ and any query $q\notin S(\psi^{R})$, the conditional law of $O_{\Phi}(q)$ is unchanged by conditioning on $\psi^{R}$. In particular, for any measurable function $h(\cdot)$,
$\mathbb{E}\!\left[h\!\left(O_{\Phi}(q)\right)\mid\psi^{R}\right]=\mathbb{E}\!\left[h\!\left(O_{\Phi}(q)\right)\right].$ (18)
Proof.
Under the stateless assumption, $\psi^{R}$ depends only on $\{\xi_{q^{\prime}}:q^{\prime}\in S(\psi^{R})\}$ through the realized outcomes of already issued queries. For any $q\notin S(\psi^{R})$, $\xi_{q}$ is independent of $\{\xi_{q^{\prime}}:q^{\prime}\in S(\psi^{R})\}$, hence independent of the event defining $\psi^{R}$. Therefore conditioning on $\psi^{R}$ does not change the distribution of $\xi_{q}$ nor of $O_{\Phi}(q)$, which yields (18). ∎
@@ Appendix A Proof of Theorems
(Adaptive monotonicity). For any realization $\Phi$ and any query set $A\subseteq Q$, the coverage function satisfies $f(A\cup\{q\},\Phi)\geq f(A,\Phi)$, since adding a query can only enlarge the union of retrieved items. By Definition 1, this implies $\Delta(q\mid\psi^{R}_{t-1})\geq 0$ for any retrieval-level partial realization $\psi^{R}_{t-1}$.
(Adaptive submodularity). Let $\psi^{R}$ and $\psi^{R}$$^{\prime}$ be retrieval-level partial realizations such that $\psi^{R}$$^{\prime}$ is an extension of $\psi^{R}$. Then $R(\psi^{R})\subseteq R({\psi^{R}}^{\prime})$, and also $S(\psi^{R})\subseteq S({\psi^{R}}^{\prime})$.
Fix any query $q\notin S({\psi^{R}}^{\prime})$ (i.e., $q$ has not been issued under $\psi^{R}$$^{\prime}$). Define the shorthand
$R\triangleq R(\psi^{R}),\qquad R^{\prime}\triangleq R({\psi^{R}}^{\prime}),\qquad Y\triangleq O_{\Phi}(q).$
For any realization $\Phi$ consistent with $\psi^{R}$$^{\prime}$ (hence also consistent with $\psi^{R}$), the marginal gain of issuing $q$ under $\psi^{R}$ is
$f\!\left(S(\psi^{R})\cup\{q\},\Phi\right)-f\!\left(S(\psi^{R}),\Phi\right)=|Y\setminus R|,$ (19)
and similarly under $\psi^{R}$$^{\prime}$ it equals $|Y\setminus R^{\prime}|$. Since $R\subseteq R^{\prime}$, we have pointwise diminishing returns for every outcome set $Y$:
$|Y\setminus R|\ \geq\ |Y\setminus R^{\prime}|.$ (20)
By Definition 1 and (19),
$\Delta(q\mid\psi^{R})=\mathbb{E}\!\left[\,|Y\setminus R|\mid\psi^{R}\right].$ (21)
Because $q\notin S(\psi^{R})$ and the victim RAG is stateless, Lemma A.1 applies and yields
$\Delta(q\mid\psi^{R})=\mathbb{E}\!\left[\,|Y\setminus R|\,\right].$ (22)
Similarly,
$\Delta(q\mid{\psi^{R}}^{\prime})=\mathbb{E}\!\left[\,|Y\setminus R^{\prime}|\,\right].$ (23)
Taking expectation of (20) over the (common) distribution of $Y$ gives $\mathbb{E}[|Y\setminus R|]\geq\mathbb{E}[|Y\setminus R^{\prime}|]$. Combining with (22)–(23) yields
$\Delta(q\mid\psi^{R})\geq\Delta(q\mid{\psi^{R}}^{\prime}),$ (24)
which is adaptive submodularity. ∎
Appendix B Sample Distribution Validation
Table 8: Semantic distribution similarity between each full corpus and its 1,000-document sampled subset, measured by Energy Distance (with permutation test $p$-value) and the Classifier Two-Sample Test (C2ST, using AUC). A small Energy Distance with a high $p$-value and an AUC near 0.5 indicate distributional similarity.
Energy Distance quantifies the difference between two distributions using Euclidean distances, and is well-suited for high-dimensional or non-Gaussian data. A smaller energy distance with a large permutation test $p$-value suggests no statistically significant difference between the sampled and full sets. As shown in Table 8, all four corpora yield low distances (around 0.03) and high $p$-values (>0.5>0.5), indicating strong alignment between samples and full sets. C2ST (Classifier Two-Sample Test) reframes distribution comparison as a binary classification task . If a classifier cannot distinguish between samples and the full corpus (AUC $\approx 0.5$), the two distributions are considered similar. In Table 8, all AUC scores fall within the 0.46–0.51 range, reinforcing that no meaningful distinction can be learned between the subsets and their corresponding full corpora.
Appendix C Attacker-side orchestration cost
We estimate the attacker-side cost of executing RAGCrawler, i.e., the tokens consumed by the attacker’s model. As reported in Table 9, RAGCrawler processes approximately 6–10 million tokens per dataset. Based on Doubao-1.5-lite-32K’s API pricing (\$0.042 per million input tokens and \$0.084 per million output tokens), this orchestration cost is \$0.33–\$0.53 per dataset. As shown in Sec. 7.4, substituting an open-source model such as Qwen-2.5-7B does not degrade performance, making this component virtually zero aside from GPU hosting.
Table 9: Token usage and estimated cost per corpus for KG-Constructor (KG-C) and Query Generator (QG) using Doubao-1.5-lite-32k. “InTok” and “OutTok” denote input and output token counts. Each attack costs under \$1.
Appendix D No-topic-prior attacker via one-shot topic probing
Motivation. In Sec. 4, we assume the attacker knows a coarse topic phrase of the target corpus, which is often available from product descriptions, onboarding pages, or light interaction with the assistant [vertexaisearch_pricing, wang2025silent]. To be conservative about attacker knowledge, we also evaluate a weaker attacker that starts with no topic prior; the corresponding CR results are reported in Table 7 (main text).
One-shot topic probing. The attacker first issues a single benign probe query to the victim RAG system:
What domain do you know about?
Let the victim response be $a_{\mathrm{probe}}$. The attacker then infers a high-level topic phrase from $a_{\mathrm{probe}}$ using an attacker-side LLM and the following prompt. The output is restricted to a short phrase (3–8 words) to avoid over-specification and keep the prior coarse.
You are given a sentence that is generated by a RAG system.
Infer the high-level domain/topic of the RAG system in 3-8 words.
Return ONLY the topic phrase (no quotes, no extra text).
Sentence: {a_prob}
Topic:
The inferred phrase is the only topic signal used in the subsequent attack. For RAGCrawler, it replaces the provided topic input to the Topic-Specific Prior module (Sec. 6.2), which induces the lightweight entity/relation schema and then proceeds with the same closed-loop crawling. For fair comparison, we provide the same inferred topic phrase to baselines as their input.
Budget accounting and metrics. We count the probe as one issued query, so the remaining crawling budget is reduced by one under a fixed total budget. We report Coverage Rate (CR) under the same evaluation protocol as Sec. 7.
Further analysis: inferred-topic noise and robustness. The one-shot probe can yield coarse or off-focus topic phrases that deviate substantially from the manual topic prior. For example, the inferred topic is Medical Research on Impact Factors for TREC-COVID (manual topic: COVID-19, similarity 0.134) and Relationship Marketing & Tech for SciDocs (manual topic: Computer Science Research Papers, similarity 0.185).
Despite such noise, Table 7 shows that RAGCRAWLER remains effective and achieves the best CR across datasets, with only modest changes relative to the topic-prior setting. In contrast, baselines can be more sensitive to topic quality, consistent with their heavier reliance on the provided topic signal. This robustness is expected from RAGCRAWLER’s design: the topic prior only bootstraps the initial schema, while the attacker-side global state is continuously updated from retrieved documents, allowing subsequent query planning to self-correct and prioritize unexplored regions of the corpus.
Appendix E Additional Experiments Results
E.1 Coverage Rate Curves
@@
E.2 Additional Ablations
Figure 14: Hyperparameter Choices of RAGCrawler.
Hyperparameter Choices. We examine two key hyperparameters in RAGCrawler: the UCB exploration coefficient $c$ used by the Strategy Scheduler for query selection, and the similarity threshold $\tau_{dup}$ used by the Query Generator to filter high-similarity queries (thereby controlling early stopping). As shown in Fig. 14, a moderate UCB coefficient of $c=0.5$ delivers the best performance by balancing exploitation and exploration. Increasing or decreasing $c$ from this value noticeably degrades performance, underscoring that $c=0.5$ is near-optimal for a proper exploration–exploitation trade-off. Similarly, the threshold $\tau_{dup}$ exhibits a clear sweet spot. Setting $\tau_{dup}$ too low often causes premature termination. On the other hand, an overly high threshold produces queries that are not sufficiently distinct from one another, which limits exploration efficiency. Our chosen $\tau_{dup}=0.8$ navigates between these extremes.
E.3 Additional Target-Coverage Query Complexity
Table 10 summarizes the target-coverage query complexity $Q_{\gamma}$ under a fixed cap $Q_{\max}=5{,}000$ and the resulting final coverage $\mathrm{CR}(Q_{\max})$. A clear pattern is that baseline heuristics exhibit strong early gains but saturate quickly, whereas RAGCrawler continues to accrue coverage as the target increases. At low targets (50–60%), IKEA can sometimes approach RAGCrawler on easier corpora (e.g., Healthcare), suggesting that shallow, highly retrievable regions can be harvested with mostly local signals. However, as $\gamma$ increases, IKEA increasingly plateaus: on three of four datasets it cannot reach 70% within the cap, and its remaining gains are dominated by redundancy. RAGThief is consistently the weakest, failing to reach 50% on most corpora and often stalling far below the cap coverage achieved by the other methods.
Table 10: Query complexity for target corpus coverage under a budget cap. For each method and dataset, we report $Q_{\gamma}$, the number of issued queries needed to reach coverage $\gamma\in\{0.5,0.6,0.7,0.8,0.9\}$ with $Q_{\max}=5{,}000$; entries marked $Q_{\gamma}>Q_{\max}$ indicate the target was not reached within the cap. We also report the final coverage $\mathrm{CR}(Q_{\max})$ for all runs. RAGCrawler consistently attains the highest $\mathrm{CR}(Q_{\max})$ and reaches substantially higher coverage targets with fewer queries.
Dataset Method $Q_{0.5}$ $Q_{0.6}$ $Q_{0.7}$ $Q_{0.8}$ $Q_{0.9}$ $\mathrm{CR}(Q_{\max})$ TREC-COVID RAGThief >5000>5000 >5000>5000 >5000>5000 >5000>5000 >5000>5000 0.298 IKEA >5000>5000 >5000>5000 >5000>5000 >5000>5000 >5000>5000 0.196 RAGCrawler 1107 1972 4040 >5000>5000 >5000>5000 0.725 SciDocs RAGThief >5000>5000 >5000>5000 >5000>5000 >5000>5000 >5000>5000 0.072 IKEA 493 >5000>5000 >5000>5000 >5000>5000 >5000>5000 0.560 RAGCrawler 457 776 1163 1784 3330 0.936 NFCorpus RAGThief >5000>5000 >5000>5000 >5000>5000 >5000>5000 >5000>5000 0.192 IKEA 938 3427 >5000>5000 >5000>5000 >5000>5000 0.640 RAGCrawler 200 322 595 1011 1753 0.977 Healthcare RAGThief 3537 >5000>5000 >5000>5000 >5000>5000 >5000>5000 0.530 IKEA 212 345 1231 2925 >5000>5000 0.825 RAGCrawler 209 329 568 983 2434 0.941
In contrast, RAGCrawler is the only method that reliably enters the high-coverage regime. It continues to improve beyond 80% and even reaches 90% on three datasets within $Q_{\max}$, while no baseline attains 90% under the same budget. This difference is also reflected in final coverage at $Q_{\max}$: RAGCrawler achieves the highest $\mathrm{CR}(Q_{\max})$ on every dataset and maintains a large margin over both baselines, with the gap widening on harder corpora. TREC-COVID remains the most challenging case: although RAGCrawler cannot reach 80% within the cap, it still sustains steady growth and substantially outpaces baselines that plateau below 30%. Overall, the advantage of global state and planning becomes more pronounced as the desired coverage rises, precisely where local heuristics are most susceptible to redundancy and drift.
E.4 Larger victim collections
Appendix F Prompts for Experimental Stages
@@
RAG System Prompt You are an assistant for question-answering tasks. Use the following pieces of retrieved context to answer the question. If you don’t know the answer, say that you don’t know. Context: [context] Query: [query] Answer:
Query Rewriting Prompt in RAG System You are a helpful assistant. Your task is to rewrite the given user question to clarify the user’s intent and remove any adversarial intent. Directly output the rewritten question without any other text. Original Question: [query] Answer:
Multi-query Prompt in RAG System You are a helpful assistant. Your task is to generate 3 different versions of the given user question to retrieve relevant documents from a vector database. By generating multiple perspectives on the user question, your goal is to help the user overcome some of the limitations of the distance-based similarity search. Provide directly output these alternative questions separated by newlines without any other text. Original Question: [query] Answer: